OpenStudy (anonymous):

Graph the function and label Zeros, Intersections, Vertices,and Asymptotes as appropriate. The function is f(x)= -2^x

OpenStudy (anonymous):

Do you know what zeroes, intersections, vertices, and asymptotes are?

OpenStudy (anonymous):

Honestly not really I'm sorry

OpenStudy (anonymous):

That's fine, I just needed to know so I know how much to explain :) Alright, a zero is when y = 0, or f(x) = 0. There shouldn't be an intersection if there's only one function. Since the function is just a single curve, there also shouldn't be a vertex. There is an asymptote at f(x) = 0. Since the asymptote is at y = 0, there are no zeroes of this function. An asymptote is a point, in this case f(x)=0, where the graph will never actually reach.

OpenStudy (anonymous):

Oh Ok Thanks!!! Really helped alot! Can you help me with one more?

OpenStudy (anonymous):

Please?

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

Alright so its the same question but the function is different. The function is f(x)= In(x+2)

OpenStudy (anonymous):

Alright, so it's similar. For these it's easy to look at a graph of the equation so I attached one here.

OpenStudy (anonymous):

so would an Asymptote be -2?

OpenStudy (anonymous):

Yep!

OpenStudy (anonymous):

Yes!! So what are the Vertices and zeros?

OpenStudy (anonymous):

There aren't vertices in graphs that have no minimum or maximum. To find the zeroes you can set f(x)=0 and solve.

OpenStudy (anonymous):

so you would set In(x+2)=0?

OpenStudy (anonymous):

@akeith15 What do the zeros end up being?

OpenStudy (anonymous):

Yes, you would set ln(x+2)=0 then solve.

OpenStudy (anonymous):

x=-1

OpenStudy (anonymous):

I knew it!! Thanks again for the help!!